OpenStudy (anonymous):

Solve x + 3y = 7 if x and y must be whole numbers.

OpenStudy (anonymous):

{(1, 2), (4, 1), (7, 0)}

OpenStudy (anonymous):

x + 3y = 7 => x = 7-3y, let y = t solution set becomes parametrized as: \[\left[\begin{matrix}x \\ y \end{matrix}\right] = \left[\begin{matrix}7-3t \\ t \end{matrix}\right]\] so let t = any positive whole number and you can find the corresponding y

OpenStudy (anonymous):

woops meant corresponding x